OpenStudy (anonymous):

Factor by grouping. 14ax-y+2ay-7x

OpenStudy (unklerhaukus):

which terms have common factors?

OpenStudy (anonymous):

first and fourth and second and third

OpenStudy (unklerhaukus):

ok so what is the common factor between: 14ax & -7x and -y & 2ay

OpenStudy (anonymous):

x, y

OpenStudy (unklerhaukus):

SO 14ax-7y = x(14a-7) and -y+2ay = y(-1+2a) is there another common factor in the first one, i.e. between 14a & -7 ?

OpenStudy (anonymous):

yes, 7

OpenStudy (unklerhaukus):

we have 14ax-y+2ay-7x = [14ax-7x]+[-y+2ay] = x[14a-7]+y[-1+2a] = 7x[2a-1]+y[-1+2a] = . . . what do we get if we switch around the terms in [-1+2a] around,

OpenStudy (anonymous):

7x(2a1-)+y(2a-1) ?

OpenStudy (unklerhaukus):

i know you mean 7x(2a-1)+y(2a-1) can you see another common factor now?

OpenStudy (anonymous):

yes that is what i meant. no i do not see another common factor

OpenStudy (unklerhaukus):

is there anything the same between 7x(2a-1) & y(2a-1) ?

OpenStudy (anonymous):

2a-1

OpenStudy (unklerhaukus):

yes, imagine (for a second) that (2a-1) = z 7xz + yz how would you factorize this?

OpenStudy (anonymous):

i really dont know

OpenStudy (unklerhaukus):

as you can see, z is the common factor, so bring it out the front: z(7x + y) now just put back z = (2a-1)

OpenStudy (anonymous):

(2a-1)(7x+y)

OpenStudy (unklerhaukus):

that is correct!

OpenStudy (anonymous):

thanks so much.

OpenStudy (unklerhaukus):

if you want to check your answer, you could foil it, and you should get the expression in the question statement

OpenStudy (anonymous):

alright great thanks